[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Aug 16, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Wed Aug 16 17:18:41 UTC 2017


Here is today's summary of NHC computer operations:
--- Monica

----------

TAFB:
Quiet day in TAFB
HSU:
Doug Gaer is working on creating a clone of the ATCF1 server on ATCF3 to  
enable a smooth back up transition from 1 to 3 should 1 fail. In order to  
do this we need approximately 2 hours during a period when no advisories  
are being issued so that Doug can take the ATCF server down for the clone.  
Since Gert may be extratropical by Friday, the proposed time (subject to  
change) is possibly Friday afternoon unless something else develops.
Other items:
1) Joe Zajic rescheduled the upgrade to the processing code for the G16  
data to early next week. TSB will update the staff again when we hear more  
from Joe.
2) Monitor upgrade work on the operations floor continues. A few more  
monitors to go.
NHC RFCs:
None scheduled for the rest of this week

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170816/763b2f6a/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list